Solve the following system by the addition method.
3x-7y= -12
3x+14y= -3
x=
Y=

Answer by ewatrrr(24785)


Hi,
Solve the following system by the addition method
3x-7y= -12
3x+14y= -3
Multiplying the first equation thru by 2 and adding the equations together
6x-14y= -24
3x+14y= -3
9x = -27
x = -3
substituting x into 3x+14y= -3
-9 + 14y = -3
14y = 6
y = 6/14 or 3/7
checking our answer with 3x-7y = -12
-9 - 3 = -12